1. Purchased equipment for 70,000.this equipment has a five year life and an 8000 residual value.calculate depreciation for each of the five years using the declining balance method at twice the straight line rate?

1
Expert's answer
2022-02-07T11:54:38-0500

Solution:

Straight-line rate = 1/5 or 20%; Declining Rate = 40%

Maximum Depreciation allowed = $62,000


"B=40\%\ of\ A
\\ C=A-B
\\ D=A-D
\\ E_n=C_n+C_{n+1}"

In Year 5, the asset may not be depreciated beyond its residual value. That is, the net

book value may not be less than the residual value. Applying the double declining

balance method in year 5 calculates an expense of (70,000 – 60,928) "\times" 40% =

$3,628.80 which reduces the book value below the residual value.